How do I find the system of linear equation for
9x+4y=-25
8x+4y=4

Answer by mananth(16946)

How do I find the system of linear equation for
9x+4y=-25...............1
8x+4y=4.................2
...
subtract 2 from 1
9x-8x=-25-4
x= -29
...
plug value of x in equation1
...
9*-29+4y=-25
-261+4y=-25
4y=236
y=236/4
y=59
(-29,59) is the solution